If a man moves a large box that weighs 10 Newton 20 meters in 30 seconds, how much power was used


1
Expert's answer
2020-10-30T13:17:45-0400

Power is the rate showing how quickly work can be done:


"P=\frac{W}{t}."


Work is force times distance:


"W=Fd."

Plug this into first equation:


"P=\frac{Fd}{t}=\frac{10\cdot20}{30}=6.7\text{ W}."